OpenStudy (anonymous):

What is the equation of the circle with center (-3.2,2.1) and radius 4.3?

OpenStudy (astrophysics):

Remember the brief explanation I showed you on your previous post? :P

OpenStudy (anonymous):

yes x2+y2=r2 would i just plug them in

OpenStudy (astrophysics):

Look at the drawing and see what I said for what the centre and radius are :)

OpenStudy (anonymous):

x+3.2^2=y-2.1^2=4.3^2

OpenStudy (anonymous):

+ not equal

OpenStudy (astrophysics):

Yeah, make sure you add brackets and such, but looks good :)

OpenStudy (anonymous):

thank uuuu xx

OpenStudy (astrophysics):

No problemo ^.^